Consider the arithmetic series 3+7+11+15+19+23+27+...and answer the questions which fo.
(a) T1= S1=
(ii) express T1 in terms of S1
(b) T2= S2=
(ii)Express T2 in terms of S2 and S1

Found 2 solutions by ikleyn, 02022256164089:
Answer by ikleyn(52788)

.

HINT

    T1 is the first term of the sequence.


    T2 is the second term of the sequence.


    Sn is the sum of the first "n" terms of the sequence.


        So, S1 is the first term of the sequence; same as T1;  therefore, S1 = T1.


            S2 = T1 + T2.   Therefore,  T2 = S2 - S1.


Happy learning (!)

---------------------

For introductory lessons on arithmetic progressions see
    - Arithmetic progressions
    - The proofs of the formulas for arithmetic progressions
    - Problems on arithmetic progressions
    - Word problems on arithmetic progressions
in this site.

Also,  you have this free of charge online textbook in ALGEBRA-II in this site
    - ALGEBRA-II - YOUR ONLINE TEXTBOOK.

The referred lessons are the part of this online textbook under the topic "Arithmetic progressions".


Save the link to this textbook together with its description

Free of charge online textbook in ALGEBRA-II
https://www.algebra.com/algebra/homework/complex/ALGEBRA-II-YOUR-ONLINE-TEXTBOOK.lesson

into your archive and use when it is needed.


Answer by 02022256164089(1)